# OTTER: A Vision-Language-Action Model with Text-Aware Visual Feature Extraction

### v3

@@ 4.1 Environment Setup
Simulation Environment We use the LIBERO benchmark (Liu et al., 2024) for simulation evaluation. Specifically, we use the tasks and datasets in LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-90, which contains diverse objects, scene layouts, and language instructions. Each simulation task has 50 demonstrations. We evaluate OTTER’s capabilities on both in-distribution tasks and unseen tasks. The in-distribution tasks are the 30 tasks in the original LIBERO-Spatial/Object/Goal, which can evaluate the model’s multi-task learning capabilities. In addition, we also construct 10 novel tasks, where we modify the language instructions and corresponding objects of 10 original tasks from LIBERO-90. For the 10 unseen tasks, we follow the same convention in LIBERO (Liu et al., 2024) about object initialization and goal configuration by defining task bddl files. Example scenes in the simulation are shown in the left column of Figure 4.
@@ 4.2 Baselines
To evaluate if the text-aware visual features extracted in OTTER can better leverage the semantic understanding capabilities of the pre-trained VLMs, we consider three baselines that directly take all visual and language features, including two state-of-the-art open-sourced VLA models and one variant of OTTER:
@@
Method Pouring Drawer Poking Pick and Place Mean±$\pm$±Std. Err.
@@
Table 2: Multi-primitive zero-shot generalization: We train models across four manipulation primitives (pouring, drawer manipulation, poking, and pick-and-place) with a total of 1,185 human tele-operated demonstration trajectories and evaluate them on 150 trials of completely unseen tasks within these primitives. Despite the inherent difficulty of zero-shot generalization across multiple primitives, OTTER achieves significantly higher success rates compared to baselines across all primitives. While baseline models struggle with generalization, particularly on pouring and drawer tasks (0% success rate), OTTER maintains substantial performance (60-93% success rate) on unseen tasks. The results further suggest that OTTER’s generalization capabilities can be enhanced through increased model capacity (OTTER-L) and pre-training on large robotic datasets (OTTER-OXE).
@@ 5.1 Real-world Experiments
Finetuned OpenVLA 85% $\pm$ 0.9%* 88% $\pm$ 0.8%* 79% $\pm$ 1.0%* 84% $\pm$ 0.9% 29% $\pm$ 0.9%
DFP-OTTER 78% $\pm$ 0.9% 80% $\pm$ 1.0% 82% $\pm$ 1.1% 80% $\pm$ 1.0% 28% $\pm$ 1.1%
OTTER (Ours) 84% $\pm$ 1.0% 89% $\pm$ 1.2% 82% $\pm$ 1.0% 85% $\pm$ 1.1% 59% $\pm$ 0.8%
Table 3: Simulation results on LIBERO. We evaluate OTTER and other baselines on 300 trials on in-distribution tasks in LIBERO-Spatial/Object/Goal, and 100 trials on unseen tasks we constructed. The numbers marked with * of are directly referred from the OpenVLA (Kim et al., 2024) paper. The detailed training and evaluation on the unseen tasks are described in Section 4.1.
@@ 5.2 Simulation Experiments
We compare OTTER with various baselines in simulation, as shown in Table 3. For fair comparison with OpenVLA and Octo, we process the simulation datasets following the procedure described in the OpenVLA paper (Kim et al., 2024), and use their reported performance on the standard LIBERO tasks. For the unseen tasks, we change 10 tasks in LIBERO-90 with different objects and distractors to test the generalizability of all the models on unseen tasks. From Table 3, we found all the models perform similarly on training tasks in LIBERO due to the limited variations of the tasks and sufficient demonstration datasets. However, for tasks in LIBERO-Object and LIBERO-Spatial, DFP-OTTER is slightly worse than OTTER because OTTER does better in localizing the object to interact with. Note that Octo and OpenVLA are pre-trained on the OXE dataset but DFP-OTTER is trained from-scratch on LIBERO, so DFP-OTTER’s performance is slightly worse. In unseen tasks, OTTER can outperform other baselines by a large margin, demonstrating its generalization capabilities to novel scenarios, which is consistent with the conclusion of the real-world experiments.
@@ Acknowledgement
Dataset usage and model training work solely conducted at Berkeley. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. This research was performed at the AUTOLAB at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS ”People and Robots” (CPAR) Initiative. In their academic roles at UC Berkeley, Letian Fu, Huang Huang, and Ken Goldberg are supported in part by donations from Meta, Autodesk, Google, Siemens, Toyota Research Institute, Bosch, and by equipment grants from PhotoNeo, Nvidia, NSF AI4OPT Centre, and Intuitive Surgical.
7 Potential Broader Impact
@@ B.1 Model Architecture for OTTER and Baselines
The details of our model parameters can be found in Table 7. All the baselines share the same hyper-parameters with OTTER. For OTTER w.o. CLIP Vision, we use a ViT Encoder based on the implementation of https://github.com/google-research/vision_transformer with a ViT-Ti/16 configuration with half of the number of attention layers. For OTTER w.o. $f_{e}$ and OTTER w.o. $f^{\prime}_{l}$, we use the same model configuration but only remove the corresponding attention pooling layers. We incorporate action chunking into OpenVLA by asking it to predict the next 16 actions, which performs better than vanilla OpenVLA which predicts only the next step. For Octo, we use the official Hugging Face Checkpoint at hf://rail-berkeley/octo-small-1.5 which is in a comparable size with our model. During inference, we cache the CLIP feature outputs. This enables the ViT-L/14 OTTER model to perform inference at $17Hz$ on a single NVIDIA 3090Ti, allowing real-time control.
@@
CLIP Model ViT-L/14
@@
# Context Length 12
# Action Prediction Horizon 12
@@
Table 7: Hyperparameters for OTTER model architecture. Values in the parenthesis shows the hyperparameters for a larger and wider OTTER.

### v4

@@ 4.1 Environment Setup
Simulation Environment We use the LIBERO benchmark (Liu et al., 2024a) for simulation evaluation. Specifically, we use the tasks and datasets in LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-90, which contains diverse objects, scene layouts, and language instructions. Each simulation task has 50 demonstrations. We evaluate OTTER’s capabilities on both in-distribution tasks and unseen tasks. The in-distribution tasks are the 30 tasks in the original LIBERO-Spatial/Object/Goal, which can evaluate the model’s multi-task learning capabilities. In addition, we also construct 10 novel tasks, where we modify the language instructions and corresponding objects of 10 original tasks from LIBERO-90. For the 10 unseen tasks, we follow the same convention in LIBERO (Liu et al., 2024a) about object initialization and goal configuration by defining task bddl files. Example scenes in the simulation are shown in the left column of Figure 4.
@@ 4.2 Baselines
To evaluate if the text-aware visual features extracted in OTTER can better leverage the semantic understanding capabilities of the pre-trained VLMs, we consider four baselines that directly take all visual and language features, including three state-of-the-art open-sourced VLA models and one variant of OTTER:
@@
[…]
@@
Table 2: Multi-primitive zero-shot generalization: We train models across four manipulation primitives (pouring, drawer manipulation, poking, and pick-and-place) with a total of 1,185 human tele-operated demonstration trajectories and evaluate them on 150 trials of completely unseen tasks within these primitives. Despite the inherent difficulty of zero-shot generalization across multiple primitives, OTTER achieves significantly higher success rates compared to baselines across all primitives. While baseline models struggle with generalization, particularly on pouring task (0% success rate), OTTER maintains substantial performance (60-93% success rate) on unseen tasks. The results further suggest that OTTER’s generalization capabilities can be enhanced through increased model capacity (OTTER-L) and pre-training on large robotic datasets (OTTER-OXE).
@@ 5.1 Real-world Experiments
Finetuned OpenVLA 85% $\pm$ 0.9%* 88% $\pm$ 0.8%* 79% $\pm$ 1.0%* 84% $\pm$ 0.9% 48% $\pm$ 1.0%
DFP-OTTER 79% $\pm$ 0.9% 80% $\pm$ 1.0% 78% $\pm$ 1.1% 79% $\pm$ 1.0% 45% $\pm$ 1.0%
OTTER (Ours) 84% $\pm$ 1.0% 89% $\pm$ 1.2% 79% $\pm$ 0.9% 84% $\pm$ 1.1% 61% $\pm$ 1.1%
Table 3: Simulation results on LIBERO. We evaluate OTTER and other baselines on 30 in-distribution tasks in LIBERO-Spatial/Object/Goal and on 10 unseen tasks we constructed, each task 50 trials. The numbers marked with * of are directly referred from the OpenVLA (Kim et al., 2024) paper. The detailed training and evaluation on the unseen tasks are described in Section 4.1.
@@ 5.2 Simulation Experiments
We compare OTTER with various baselines in simulation, as shown in Table 3. For fair comparison with OpenVLA and Octo, we process the simulation datasets following the procedure described in the OpenVLA paper (Kim et al., 2024), and use their reported performance on the standard LIBERO tasks. For the unseen tasks, we change 10 tasks in LIBERO-90 with different objects and distractors to test the generalizability of all the models on unseen tasks. From Table 3, we found all the models perform similarly on training tasks in LIBERO due to the limited variations of the tasks and sufficient demonstration datasets. For tasks in LIBERO-Object and LIBERO-Spatial, DFP-OTTER is worse than OTTER because OTTER does better in localizing the object to interact with. Note that Octo and OpenVLA are pre-trained on the OXE dataset but DFP-OTTER is trained from-scratch on LIBERO, so DFP-OTTER’s performance is worse. In unseen tasks, OTTER can outperform other baselines by a large margin, demonstrating its generalization capabilities to novel scenarios, which is consistent with the conclusion of the real-world experiments.
@@ 5.4 Scaling up Vision and Language Encoders
5.4.1 Scaling up Pre-training Dataset with Human Videos
Dataset Prior work like Octo and OpenVLA has studied pre-training on large robotic datasets like Open X-Embodiment dataset. To leverage a broader source of data such as human videos, we construct a human dataset consisting of the DexYCB dataset (Chao et al., 2021) and the HOI4D dataset (Liu et al., 2024b). Both datasets provide labeled human hand poses in the camera frame. We post-processed the datasets to extract human hand motion trajectories. In total, we obtain 15.6k human hand motion trajectories consisting of RGB image frames, language instructions and human hand actions from the DexYCB dataset and the HOI4D dataset.
Results We first pre-train OTTER using human videos, then finetune it using our robot dataset. We evaluate the benefits of pre-training on this human video dataset. Results are shown in Table 5. The human dataset has different camera angles and no wrist view image and the proprioception of the human dataset is also in various coordinate frames. Despite the significant discrepancy between the human and robot dataset, we found naively pre-training OTTER on human dataset shows improved performance on the unseen tasks, indicating that pre-traing on human videos can improve the generalization ability of the model. More details and discussions about the human dataset and training regimes can be found in Appendix C.1.
Method Training Tasks Unseen Tasks
OTTER w.o. Human Pre-training 68% $\pm$ 4.3% 62% $\pm$ 4.2%
OTTER w. Human Pre-training 67% $\pm$ 4.0% 70% $\pm$ 4.5%
Table 5: Physical results of OTTER with and without human pre-training evaluated on 100 trials for in-distribution training tasks and 70 trials for unseen tasks. Pre-training OTTER on human dataset improves the performance on unseen tasks.
@@ Acknowledgement
Dataset usage and model training work solely conducted at Berkeley. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. This research was performed at the AUTOLAB at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS ”People and Robots” (CPAR) Initiative. In their academic roles at UC Berkeley, Letian Fu, Fangchen Liu, Huang Huang, and Ken Goldberg are supported in part by donations from Meta, Autodesk, Google, Siemens, Toyota Research Institute, Bosch, and by equipment grants from PhotoNeo, Nvidia, NSF AI4OPT Centre, and Intuitive Surgical.
Impact Statement
@@ B.1 Model Architecture for OTTER and Baselines
The details of our model parameters can be found in Table 8. All the baselines share the same hyper-parameters with OTTER. For OTTER w.o. CLIP Vision, we use a ViT Encoder based on the implementation of https://github.com/google-research/vision_transformer with a ViT-Ti/16 configuration with half of the number of attention layers. For OTTER w.o. $f_{e}$ and OTTER w.o. $f^{\prime}_{l}$, we use the same model configuration but only remove the corresponding attention pooling layers. We incorporate action chunking into OpenVLA by asking it to predict the next 16 actions, which performs better than vanilla OpenVLA which predicts only the next step. For Octo, we use the official Hugging Face Checkpoint at hf://rail-berkeley/octo-small-1.5 which is in a comparable size with our model. During inference, we cache the CLIP feature outputs. This enables the ViT-L/14 OTTER model to perform inference at $50Hz$ on a single NVIDIA 3090Ti, allowing real-time control.
@@
CLIP Model (Real) ViT-L/14
CLIP Model (Sim) ViT-L/32
@@
# Context Length (Real) 12
# Context Length (Sim) 4
# Action Prediction Horizon (Real) 12
# Action Prediction Horizon (Sim) 1
@@
Table 8: Hyperparameters for OTTER model architecture. Values in the parenthesis shows the hyperparameters for a larger and wider OTTER for the real world experiments.
@@ Appendix C Vision-Language Attention Visualization
C.1 Pre-training on the Human Dataste
Both datasets provide labeled human hand poses in the camera frame, represented by the 3d positions of the 21 MANO (Romero et al., 2017) hand joints. We post-processed the labeled hand pose to extract the human hand wrist transformation as the human state, where the translation is the human wrist joint position and the rotation is the palm rotation, estimated by fitting the palm to a plane from the metacarpophalangeal joint locations of the five fingers and the wrist joint. DexYCB dataset is a dataset consisting third-person view videos of a human grasps different objects from a tabletop environments. We label each image frame with the language instruction of ”pick up the {object name}”, where {object name} is the object being grasped by the human. HOI4D dataset is a dataset consisting of egocentric view videos of a human manipulating different objects. Different manipulation tasks are considered such as ”pick up objects”, ”push objects” and ”open and close the door”. We use the provided language instructions to label each image frame. In total, we obtain 8k trajectories of RGB image frames, language instructions and human states from the DexYCB dataset and 7.6k trajectories from the HOI4D dataset, resulting 15.6k human trajectories.